OpenStudy (anonymous):

A boat costs $18,550 and decreases in value by 9% per year. How much will the boat be worth after 8 years?

OpenStudy (johnweldon1993):

Value = Cost(initially)(1-r)^t \[V = C(1-r)^t\] You have 'C' = 18550 you have 'r' = .09 and you have 't' = 8 years plug those in and solve for 'V'

OpenStudy (anonymous):

thanks man i appreciate it !

OpenStudy (johnweldon1993):

no problem!